OpenStudy (anonymous):

You have a large bag in front of you and randomly pick out one candy. Find P(orange), the probability that you select a orange candy. Orange: ?, Yellow: 0.30, Brown: 0.40

OpenStudy (anonymous):

turns out the answer was 0.30

OpenStudy (anonymous):

i have another one to

OpenStudy (anonymous):

I can try to answer it

OpenStudy (anonymous):

Find P(X>2). 1: 0.30, 2: 0.40. 3: 0.15, 4: ?

OpenStudy (anonymous):

Ohhh...um.....yeah I don't know how to do that sorry...

OpenStudy (queelius):

Which events are greater than two? Then, sum them up. P[X=3 or X=4].

OpenStudy (anonymous):

would it be 0.30?

OpenStudy (queelius):

The reason the previous answer was 0.3 is this: if the events are mutually exclusive and exhaustive, then realize that the probability of *something* happening (one of the events) is necessarily 1. Then, P[orange] + P[yellow] + P[brown] = 1.

OpenStudy (queelius):

Yes, P[X=4] must be 0.15. Thus, since these events are mutually exclusive, P[X=3 or X=4] = P[X=3] + P[X=4].

OpenStudy (queelius):

Which is, as you indicated, 0.15+0.15 = 0.3.

OpenStudy (anonymous):

your a life saver. thanks so much

OpenStudy (queelius):

Another way to answer that question: P[X>2] = 1-P[X<=2]

OpenStudy (queelius):

P[X<=2] -> P[X=1 or X=2] = P[X=1] + P[X=2] (since they are mutually exclusive events)

OpenStudy (anonymous):

what if 1: 0.30, 2: 0.40, 3: 0.15, 4: 0.15 and i have to find P( not a 3)

OpenStudy (queelius):

P[X is not 3] = 1 - P[X=3] Or, alternatively: P[X is not 3] = P[X=1] + P[X=2] + P[X=4]

OpenStudy (anonymous):

so... the answer is?